Kathy needs to replace the 1000-ft2 heat exchanger she purchased 6 years ago with a 3000-ft2 heat exchanger . If she purchased the 1000-ft2 heat exchanger for $5,000 6 years ago, estimate the cost of the 3000-ft2 heat exchanger today if the power sizing exponent is 0.6 and the cost index for compressors 6 years ago was 150 and today it is 210.
Answer #1

Initial Price index of compressors =PIo=150

Price index of compressors now=PI1=210

Initial price of 1000-ft2 compressor=Co=$5000

Current price of 1000-ft2 compressor=Co*(PI1/PIo)=5000*(210/150)=$7000

Power sizing component=x=0.6

Current estimated price of 3000 ft2 compressor now=Co*(Required capacity/initial capacity)x

Current estimated price of 3000 ft2 compressor now=7000*(3000/1000)^0.6=$13532.27